A manager buys three shares of stock today, and then sells one of those shares each year for the next 3 years. His actions and the price history of the stock are summarized below. The stock pays no dividends.

Time Price Action
0 $ 150 Buy 3 shares
1 180 Sell 1 share
2 180 Sell 1 share
3 180 Sell 1 share

a. Calculate the time-weighted geometric average return on this portfolio. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

b. Calculate the time-weighted arithmetic average return on this portfolio. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

c. Calculate the dollar-weighted average return on this portfolio. (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

a.time-weighted geometric average return (gn)

gn = (1+rc)1/n-1 where rc = cumulative return over the period; n = number of periods; 1/n = 1/3= 0.33

rate of return for year 1 = (180 x 3) - (150 x 3) / (150 x 3) = 20%

rate of return for year 2 = (180 x 2) - (150 x 2) / (150 x 2) = 20%, one share is sold at the end of year 1

rate of return for year 3 = (180 x 1) - (150 x 1) / (150 x 1) = 20%, two shared sold at the end of year 2.

rc = 20% + 20% + 20% = 60% or 0.60

gn = (1 + 0.60)1/3 - 1 = 1.17 - 1 = 0.17 or 17%.

b.time-weighted arithmetic average return (an)

an = (r1 + r2 + r3) / n where rn is the rate of return for the respective year.

since the rate of return for each year is already calculated above, arithmetic mean is as below:

an = (20%+20%+20%) / 3 = 20% or 0.20.

c.dollar-weighted arithmetic average return (dn)

Year Return (r) Weight (w) r x w
1 90 1 90
2 60 2 120
3 30 3 90
Total 6 300
Weighted average 50
Return % 33%